# Minimal Subsampled Rank-1 Lattices for Multivariate Approximation with Optimal Convergence Rate

### v2

@@ 7 Numerical results
  • the least squares approximation (4.1) using the full and subsampled lattice where the frequency index set =$\mathcal{B}=\{\bm{h}\in\mathds{Z}^{d}:r_{d,\alpha,\bm{\gamma}}(\bm{h})\leq M\}$ and $M=1/(2\sqrt{\mathcal{S}_{n}(\bm{z})})$ is chosen as in Theorem 5.3, and
@@
With $n$ being the next biggest prime from a power of 2, we compute the generating vector for a lattice using the CBC algorithm from [GS25] rather than the original algorithm from [CKNS20]. The new algorithm has the advantages that it is much simpler to implement, is extensible to any dimension $d$, and (especially for SPOD weights) is cheaper to run. Empirically it gives similar $L_{2}$-approximation error bounds, but equivalent theoretical error bounds are yet to be proven. For the methods using subsampled points, we generate a random subset of size $|J|=\lceil\sqrt{n}\,\log n\rceil$, which is of the same order as in Theorem 5.3 but ignores the involved constants. The $L_{2}$-error is then estimated using $50$ random shifts of the initial lattice,
@@ 7.1 Kink function
$f(\bm{x})=\Big{(}\frac{5^{3/4}15}{4\sqrt{3}}\Big{)}^{d}\prod_{j=1}^{d}\max\Big{\{}0,\frac{1}{5}-\Big{(}x_{j}-\frac{1}{2}\Big{)}^{2}\Big{\}}$ (7.1)
in dimensions $d=2$ and $d=5$. The 2-dimensional example is depicted in Figure 2 (left).
$(0,0)$$(0,1)$$(1,1)$Refer to caption012
$(0,0)$$(0,1)$$(1,1)$Refer to caption12
Figure 2: “Kink” function (7.1) (left) and “reciprocal” function (7.2) with $q=2.5$ (right), both in dimension $d=2$.
This function was considered in [BKUV17, KUV19, BKPU22] and we know that $f\in H_{d,3/2-\varepsilon,\bm{\gamma}}(\mathds{T}^{d})$ for $\varepsilon>0$ and $\gamma_{\mathfrak{u}}=1$ for all $\mathfrak{u}\subseteq\{1,\dots,d\}$.
@@
The result for dimensions $d=2$ and $d=5$ are depicted in Figure 3.
@@
Figure 3: $L_{2}$-error of the approximations of the kink function, with $d=2$ (top) and $d=5$ (bottom) with respect to the number of sampling points for the classical lattice algorithm, the least squares method, and the kernel method using the full and subsampled lattice.
We start by evaluating the 2-dimensional example, hoping to observe the theoretical convergence rate. All rates displayed in the legend were estimated by fitting the last ten data points.
  • $\bullet$:
    For $(n,\sqrt[4]{\mathcal{S}_{n}(\bm{z})})$ we expect a decay rate arbitrarily close to $1/2$ according to Theorem 5.1. The numerics yield a rate of $0.47$ affirming this behavior.
[…]
    The classical lattice algorithm and the least squares method using the full lattice differ only by a factor of two in the radius of their frequencies. According to Theorem 2.1 they behave at least as well as $(n,\sqrt[4]{\mathcal{S}_{n}(\bm{z})})$ with a rate of $1/2$. In fact, the same rate of $0.7$ is observed in both methods and they only differ in the corresponding constant. The better than expected rate can be explained by these methods “picking up” the smoothness. Since the smoothness does not affect the shape of the frequency index set, the method for higher smoothness is the same. In our case we have smoothness $3/2-\varepsilon$, of which half is “picked up”. That it is only half is intuitively explained by full lattices only achieving half the optimal order of convergence, cf. [BKUV17].
  • $\blacktriangle$:
    For the kernel method using the full lattice we know by the discussion at the end of Section 2 that it performs at least as well as the classical lattice algorithm with a rate of $1/2$ in the worst-case setting. It was further shown in [SK25] that the kernel method (using any set of points) doubles the rate of convergence when the function is doubly as smooth. In our case we do not have quite double the smoothness, but still see this effect numerically with a rate of $1.01$.
    Note, that the experiment is only about the approximation error of one individual function which could perform better than the worst-case error.
  • $\circ$:
    For the subsampled methods we expect to achieve the same worst-case error but with fewer points. With $(|J|/\log|J|,\sqrt[4]{\mathcal{S}_{n}(\bm{z})})$, we observe a rate of $0.93$, which is about double of $0.47$ from the full lattice $(n,\sqrt[4]{\mathcal{S}_{n}(\bm{z})})$ ( $\bullet$ ).
  • $\square$ :
    According to Theorem 4.7 the least squares approximation using the subsampled lattice has at least the same rate as $(|J|/\log|J|,\sqrt[4]{\mathcal{S}_{n}(\bm{z})})$ ( $\circ$ ), which was $0.93$. For reasons explained above, the least squares method is able to “pick up” the smoothness, which is $3/2-\varepsilon$. Numerically we observe a rate of $1.37$.
  • $\triangle$ :
@@ 7.2 Reciprocal function
$f(\bm{x})=\frac{1}{a(\bm{x})}\,,\quad a(\bm{x})=1+0.5\sum_{j=1}^{d}j^{-q}\sin(2\pi x_{j})\,,$ (7.2)
with decay parameter $q=6$ and $q=2.5$. The function in two dimensions is depicted in Figure 2 (right). This function was considered in e.g., [KKNS25, Section 4.2] and solves the algebraic equation $a(\bm{x})f(\bm{x})=1$, mimicking the features of a partial differential equation with a random coefficient whilst avoiding the complexity of a spatial variable or the need of a finite element solver.
@@
Following the same procedure as in Section 7.1, in Figure 4 we depict the results for $q=6$, with $\alpha=1$ (top) and $\alpha=2$ (bottom).
@@
Figure 4: $L_{2}$-error of the approximations of the reciprocal function in $d=100$ dimensions and decay parameter $q=6$, with smoothness parameter $\alpha=1$ (top) and $\alpha=2$ (bottom), with respect to the number of sampling points for the classical lattice algorithm, the least squares method, and the kernel method using the full and subsampled lattice.
@@
Figure 5: $L_{2}$-error of the approximations of the reciprocal function in $d=100$ dimensions and with decay parameter $q=2.5$ and smoothness parameter $\alpha=2$ with respect to the number of sampling points for the classical lattice algorithm, the least squares method, and the kernel method using the full and subsampled lattice.
@@
In Figure 4 we plot the corresponding results with the relatively slow decay parameter $q=2.5$ and $\alpha=2$. This problem is considerably harder. The empirical rates drop significantly to be just above 1 for both the kernel method and least squares approximation with the subsampled lattice.
The frequencies used in the $q=6$ example in Figure 4 by the least squares approximation were nonzero only in the first $11$ and $23$ components for $\alpha=1$ and $\alpha=2$, respectively, i.e., the least squares method treated the remaining dimensions as irrelevant, while still achieving errors of $10^{-6}$ and $10^{-7}$. For the $q=2.5$ case in Figure 5 frequencies with nonzero components were encountered only in the first $52$ dimensions. This indicates that the “effective dimension” is low for the $q=6$ example compared to the $q=2.5$ example. In this case all the methods for the $q=2.5$ example behave a lot more like $(n,\mathcal{S}_{n}(\bm{z}))$ for the full lattice and $(|J|,\mathcal{S}_{n}(\bm{z}))$ for the subsampled lattice, in comparison to the $q=6$ example. It is also an example in which the subsampled least squares error shows a slightly faster empirical rate of $1.07$ than the subsampled kernel method empirical rate $1.02$.
7.3 Computational complexity
Since we are interested in the computational complexity, we measure the computation time and counted the number of iterations for the presented methods for approximating the reciprocal function with $q=2.5$ and $\alpha=2$. Both the LSQR method for the least squares approximation and the CG method for the kernel method were equipped with a stopping criterion of the relative error being smaller than $10^{-8}$. The iteration count of the subsampled methods is depicted in Figure 6.
@@ 7.3 Computational complexity
Figure 6: Number of iterations with respect to the number of points for subsampled least squares approximation and kernel method.
While the least squares approximation never exceeded $16$ iterations, the kernel method used up to $471$ iterations, growing like $|J|^{0.57}$. This affirms the predicted cost with respect to the number of iterations in Table 1.
For the actual cost, including precomputation and the matrix-vector products, we measure the elapsed time. For the subsampled lattice we used the naive matrix-vector implementation as well as the FFT implementation discussed in Section 6. Further, we modeled two scenarios: function evaluations to have no cost to evaluate and second, needing $0.1$ seconds to evaluate. The latter resembles practical examples with function evaluations corresponding to PDE solutions or similar. The results are depicted in Figure 7.
$10^{-4}$$10^{-3}$$10^{-2}$$10^{-1}$$10^{0}$$10^{-2}$$10^{0}$$10^{2}$$10^{-4}$$10^{-3}$$10^{-2}$$10^{-1}$$10^{0}$$10^{0}$$10^{2}$$10^{4}$efer to caption
@@
Figure 7: $L_{2}$-error of the numerical experiment with respect to the computation time. Left: function evaluation costs are not accounted for. Right: cost of single function evaluation is $0.1$ seconds.
@@
With the addition of a small cost of $0.1$ seconds for function evaluations, the graph resembles more the sampling complexity cost in Figure 5. Thus, in the right-hand setting when function evaluations are costly, the overhead of using an iterative solver with the the subsampling techniques of this paper is justified. If the function in question is smooth, the kernel method excels due to the doubling of the rate effect [SK25].
8 Concluding remarks
In this paper we used random subsampling to regain the optimal polynomial order of convergence for lattices while still maintaining the structure to apply fast Fourier algorithms. The focus is on having a small initial lattice size, which determines the computational complexity of the approximation algorithms. We used the least squares approximation and the kernel method. For Korobov spaces we achieve the optimal polynomial order of convergence $(|J|/\log|J|)^{-\alpha+\varepsilon}$ for $0<\varepsilon<\alpha$, with the initial lattice size $n\lesssim(|J|/\log|J|)^{2/\sqrt{1-\varepsilon/\alpha}}$, which is arbitrarily close to quadratic in terms of the actual subsampled point set size $|J|$. This improves on earlier results in [BKPU22], where there is a polynomial gap for the initial lattice size.
@@ 8 Concluding remarks
Acknowledgement. We acknowledge the financial support from the Australian Research Council Discovery Project DP240100769. This research includes computations using the computational cluster Katana supported by Research Technology Services at UNSW Sydney.
[…]

### v3

@@ 7. Numerical results
  • the least squares approximation (4.1) using the full and subsampled lattice where the frequency index set $\mathcal{B}=\{\bm{h}\in\mathds{Z}^{d}:r_{d,\alpha,\bm{\gamma}}(\bm{h})\leq M\}$ and $M=1/(2\sqrt{\mathcal{S}_{n}(\bm{z})})$ is chosen as in Theorem˜5.3, and
@@
With $n$ being the next biggest prime from a power of 2, we compute the generating vector for a lattice using the CBC algorithm from [GS25] rather than the original algorithm from [CKNS20]. The new algorithm has the advantages that it is much simpler to implement, is extensible to any dimension $d$, and (especially for SPOD weights) is cheaper to run. Empirically it gives similar $L_{2}$-approximation error bounds, but equivalent theoretical error bounds are yet to be proven. For the methods using subsampled points, we generate a random subset of size $|J|=\lceil\sqrt{n}\,\log n\rceil$, which is of the same order as in Theorem˜5.3 but ignores the involved constants. The $L_{2}$-error is then estimated using $50$ random shifts of the initial lattice,
@@ 7.1. Kink function
(7.1) $f(\bm{x})=\Big(\frac{5^{3/4}15}{4\sqrt{3}}\Big)^{d}\prod_{j=1}^{d}\max\Big\{0,\frac{1}{5}-\Big(x_{j}-\frac{1}{2}\Big)^{2}\Big\}$
in dimensions $d=2$ and $d=5$. The 2-dimensional example is depicted in Figure˜2 (left). This function was considered in [BKUV17, KUV19, BKPU22] and we know that $f\in H_{d,3/2-\varepsilon,\bm{\gamma}}(\mathds{T}^{d})$ for $\varepsilon>0$ and $\gamma_{\mathfrak{u}}=1$ for all $\mathfrak{u}\subseteq\{1,\dots,d\}$.
(0,0)(0,0)(0,1)(0,1)(1,1)(1,1)Refer to caption012
(0,0)(0,0)(0,1)(0,1)(1,1)(1,1)Refer to caption12
Figure 2. “Kink” function (7.1) (left) and “reciprocal” function (7.2) with $q=2.5$ (right), both in dimension $d=2$.
@@
The result for dimensions $d=2$ and $d=5$ are depicted in Figure˜3.
We start by evaluating the 2-dimensional example, hoping to observe the theoretical convergence rate. All rates displayed in the legend were estimated by fitting the last ten data points. For the methods using subsampled lattices, we have seen in Corollary˜5.4 that the logarithm is negligible in the asymptotic regime, but as every computation will be in the preasymptotic regime, we opted for fitting the rate to $|J|/\log|J|$ instead of $|J|$ itself, which is closer to Theorem˜5.3.
  • $\bullet$:
[…]
  • $\blacktriangle$:
    For the kernel method using the full lattice we know by the discussion at the end of Section˜2 that it performs at least as well as the classical lattice algorithm with a rate of $1/2$ in the worst-case setting. It was further shown in [SK25] that the kernel method (using any set of points) doubles the rate of convergence when the function is doubly as smooth. In our case we do not have quite double the smoothness, but still see this effect numerically with a rate of $1.01$. Note, that the experiment is only about the approximation error of one individual function which could perform better than the worst-case error.
  • $\circ$:
[…]
  • $\square$ :
[…]
  • $\triangle$ :
@@
Figure 3. $L_{2}$-error of the approximations of the kink function, with $d=2$ (top) and $d=5$ (bottom) with respect to the number of sampling points for the classical lattice algorithm, the least squares method, and the kernel method using the full and subsampled lattice.
@@ 7.2. Reciprocal function
(7.2) $f(\bm{x})=\frac{1}{a(\bm{x})}\,,\quad a(\bm{x})=1+0.5\sum_{j=1}^{d}j^{-q}\sin(2\pi x_{j})\,$,
with decay parameter $q=6$ and $q=2.5$. The function in two dimensions is depicted in Figure˜2 (right). This function was considered in e.g., [KKNS25, Section 4.2] and solves the algebraic equation $a(\bm{x})f(\bm{x})=1$, mimicking the features of a partial differential equation with a random coefficient whilst avoiding the complexity of a spatial variable or the need of a finite element solver.
@@
Following the same procedure as in Section˜7.1, in Figure˜4 we depict the results for $q=6$, with $\alpha=1$ (top) and $\alpha=2$ (bottom).
@@
Figure 4. $L_{2}$-error of the approximations of the reciprocal function in $d=100$ dimensions and decay parameter $q=6$, with smoothness parameter $\alpha=1$ (top) and $\alpha=2$ (bottom), with respect to the number of sampling points for the classical lattice algorithm, the least squares method, and the kernel method using the full and subsampled lattice.
@@
Figure 5. $L_{2}$-error of the approximations of the reciprocal function in $d=100$ dimensions and with decay parameter $q=2.5$ and smoothness parameter $\alpha=2$ with respect to the number of sampling points for the classical lattice algorithm, the least squares method, and the kernel method using the full and subsampled lattice.
@@
In Figure˜4 we plot the corresponding results with the relatively slow decay parameter $q=2.5$ and $\alpha=2$. This problem is considerably harder. The empirical rates drop significantly to be just above 1 for both the kernel method and least squares approximation with the subsampled lattice.
The frequencies used in the $q=6$ example in Figure˜4 by the least squares approximation were nonzero only in the first $11$ and $23$ components for $\alpha=1$ and $\alpha=2$, respectively, i.e., the least squares method treated the remaining dimensions as irrelevant, while still achieving errors of $10^{-6}$ and $10^{-7}$. For the $q=2.5$ case in Figure˜5 frequencies with nonzero components were encountered only in the first $52$ dimensions. This indicates that the “effective dimension” is low for the $q=6$ example compared to the $q=2.5$ example. In this case all the methods for the $q=2.5$ example behave a lot more like $(n,\mathcal{S}_{n}(\bm{z}))$ for the full lattice and $(|J|,\mathcal{S}_{n}(\bm{z}))$ for the subsampled lattice, in comparison to the $q=6$ example. It is also an example in which the subsampled least squares error shows a slightly faster empirical rate of $1.07$ than the subsampled kernel method empirical rate $1.02$.
7.3. Computational complexity
Since we are interested in the computational complexity, we measure the computation time and counted the number of iterations for the presented methods for approximating the reciprocal function with $q=2.5$ and $\alpha=2$. Both the LSQR method for the least squares approximation and the CG method for the kernel method were equipped with a stopping criterion of the relative error being smaller than $10^{-8}$. The iteration count of the subsampled methods is depicted in Figure˜6.
@@ 7.3. Computational complexity
Figure 6. Number of iterations with respect to the number of points for subsampled least squares approximation and kernel method.
While the least squares approximation never exceeded $16$ iterations, the kernel method used up to $471$ iterations, growing like $|J|^{0.57}$. This affirms the predicted cost with respect to the number of iterations in Table˜1.
For the actual cost, including precomputation and the matrix-vector products, we measure the elapsed time. For the subsampled lattice we used the naive matrix-vector implementation as well as the FFT implementation discussed in Section˜6. Further, we modeled two scenarios: function evaluations to have no cost to evaluate and second, needing $0.1$ seconds to evaluate. The latter resembles practical examples with function evaluations corresponding to PDE solutions or similar. The results are depicted in Figure˜7.
$10^{-4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$^{-2}1001$$^{0}$$10^{2}$$10^{-4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{0}$$10^{2}$$10^{4}$Refer to caption
@@
Figure 7. $L_{2}$-error of the numerical experiment with respect to the computation time. Left: function evaluation costs are not accounted for. Right: cost of single function evaluation is $0.1$ seconds.
@@
With the addition of a small cost of $0.1$ seconds for function evaluations, the graph resembles more the sampling complexity cost in Figure˜5. Thus, in the right-hand setting when function evaluations are costly, the overhead of using an iterative solver with the the subsampling techniques of this paper is justified. If the function in question is smooth, the kernel method excels due to the doubling of the rate effect [SK25].
8. Concluding remarks
In this paper we used random subsampling to regain the optimal polynomial order of convergence for lattices while still maintaining the structure to apply fast Fourier algorithms. The focus is on having a small initial lattice size, which determines the computational complexity of the approximation algorithms. We used the least squares approximation and the kernel method. For Korobov spaces we achieve the optimal polynomial order of convergence $|J|^{-\alpha+\varepsilon}$ for $0<\varepsilon<\alpha$, with the initial lattice size $n\lesssim|J|^{2/\sqrt{1-\varepsilon/\alpha}}$, which is arbitrarily close to quadratic in terms of the actual subsampled point set size $|J|$. This improves on earlier results in [BKPU22], where there is a polynomial gap for the initial lattice size.
@@ 8. Concluding remarks
Acknowledgement. T authors thank Dirk Nuyens and Daniel Potts for insightful discussions and valuable suggestions during the preparation of this work. We also thank both anonymous referees for their careful analysis and support. Further, we acknowledge the financial support from the Australian Research Council Discovery Project DP240100769. This research includes computations using the computational cluster Katana supported by Research Technology Services at UNSW Sydney.